Eric Beitz March 2000

TEXshade: frequently asked questions

This is the first release of a FAQ list for TEXshade. Feel free to contact me if you have problems, questions or suggestions about the package. I will post them and provide hopefully helpful hints in future issues of this list.

email: eric.beitz@uni-tuebingen.de
TEXshade: www.homepages.uni-tuebingen.de/beitz/tse.html

A. Problems using TEXshade

1. I cannot TEX the manual because I get the error message ‘!

TeX capacity exceeded, sorry ...’.

TEXshade needs a lot of memory for setting and shading alignments. The manual is a good test for your memory settings because it uses many alignments and fingerprints, which are in particular memory consuming. If you do not know how to increase TEX’s memory settings, and you do not know a TEX wizard either, then visit the TEXshade home-

page at http://www.homepages.uni-tuebingen.de/beitz/tse.html

to download the manual in either of three formats: DVI, PDF or PostScript.

2. I can set my alignment only when I reduce the number of basepairs by about 11,000. Otherwise I get the ‘! TeX capacity

exceeded, sorry ...’ error.

There are several parameters defining TEX’s usable space. If you are a TEX wizard (or you know one) increase the values that TEXshade complains about during the run in order to set bigger alignments. But do not be disappointed when your TEX system will not set an alignment containing thousands of residues. There is definitely an upper limit (probably the new L^{A}TEX3 will allow you to use even more memory). Setting alignments is a big job for a typesetting system!

3. I want to align 80 sequences but I get the ‘! No room for a new count’ message.

For each sequence two counter variables are used by TEXshade, further 14 counters for other purposes are needed (and TEX can handle only 255 counters). This limits the amount of sequences to about 100 in theory. But L^{A}TEX itself and each of the loaded packages allocates more counters further reducing the maximum number of sequences.

4. I receive error messages ‘! Missing $ inserted’ when TEXing my alignment. What is wrong?

At least one of the sequence names in the alignment file contains an underscore ‘ ’ symbol. This makes TEX to believe you missed to enter math mode because subscript initiated by an underscore is only allowed in math. You need to change the sequence name(s) either in the alignment file using the ‘find & replace’ option of your editor or by using the \nameseq command in the TEXshade environment. Nevertheless, subscript and superscript are permitted in sequence names,

e. g. \nameseq{1}{Name$_{sub}^{super}$} will result in Name_{sub}^{super}.

5. My sequence names start out with a number in the alignment file. Why are they ignored by TEXshade?

TEXshade analyzes the first character of each line in the alignment file in order to decide whether it is a comment, a ruler or a sequence line etc. All lines starting out with a non-letter character are interpreted as nonsequence lines. Hence, you have to change those names in the alignment file. If you want to have sequence names starting with a number you can use the \nameseq command in the TEXshade environment to introduce

the number, e. g. \nameseq{1}{57th sequence}.

6. There is an incompatiblity between TEXshade (v1.2) and the multi-language package ‘babel’!

You are right! The command \language is defined in both packages which leads to error messages. This bug is fixed since the release of TEXshade version 1.3 from March 2000. In this version \language is re-

placed by two commands: \germanlanguage and \englishlanguage.

B. Changing the output

1. How can I force TEXshade to print more residues per line?

Use the \residuesperline* command with the ‘*’ extension. This will allow you to set any number of residues per line that is desired, e. g.

\residuesperline*{97}. But then expect numerous ‘! Overfull

hbox’ errors due to printing lines that are broader than the preset \textwidth. The same command without the ‘*’ will calculate the highest number of residues fitting in one line and round it to be divisible by five.

2. Is it possible to add a caption to the TEXshade output?

Yes, it is. You can embed the TEXshade environment into a ‘figure’ with a \caption. The only thing you have to keep in mind is, that TEX cannot handle a figure environment that is bigger than one page. Thus, as long as your sequence alignment fits on one page and there is still enough space for the caption, it works fine. Otherwise you get ‘! Caption out of float’ messages. Nevertheless, in order to add a caption to an alignment that is bigger than one page you have the option to split the alignment into two parts and add the caption to one of them. Further, you could try to append an ‘empty’ figure with a caption to the alignment.

3. My alignment file contains the letters ‘B’ and ‘Z’ for Asx and Glx, respectively. How can I apply a special shading for these?

Use \funcgroup to define ‘B’ and ‘Z’ as functional groups and assign the colors and the printing style, e. g.

\funcgroup{B}{White}{Blue}{upper}{up}
\funcgroup{Z}{White}{Red}{upper}{up}

or add the new residues to an existing group, e. g.

\funcgroup{acidic/amide}{DENQBZ}{Black}{Green}{upper}{up}.

4. How can I build a legend using the ‘shadebox’ command?

The \shadebox command simply prints a color-filled box at the very location it occurs in the text. This means you have to use \shadebox in the normal text rather than inside the TEXshade environment. You find a minimal example below:

\begin{texshade}{alignmentfile.MSF}

further commands, if needed

\end{texshade}

Legend:

    \shadebox{conserved}: conserved residues
    \shadebox{White}: boring residues

\shadebox{Red}: exciting residues